What is the volume in liters occupied by 1.29 g of Freon-12 gas, CCl2F2, at 1.60 atm and 35°C?
Answer #1

No. of moles of Freon = (1.29 g) / (120.91 g/mol) = 0.0107 mol

Pressure (P) = 1.60 atm

Temperature (T) = 35 oC = 308 K

Gas constant (R) = 0.0821 L atm mol-1 K-1

PV = nRT

Volume of freon (V) = nRT / P = (0.0107 mol x 0.0821 x 308 K) / (1.60 atm) = 0.17 Liters